OpenStudy (anonymous):

find value k making h(x) continuous on [0,5]: h(x)= e^kx (0 is less than or equal to x is less than 2) x+11 (2 is less than or equal to x is less than or equal to 5)

zepdrix (zepdrix):

\[\Large\rm h(x)=\cases{e^{kx}, &$0\le x\le2$\\ \rm x+11, &$2\le x\le5$}\]So this is our function huh? :d

OpenStudy (anonymous):

yes! (sorry for the late reply)

zepdrix (zepdrix):

For continuity, we'll want the pieces to connect at x=2. We look at the limit from the left and right sides of the function.

zepdrix (zepdrix):

\[\Large\rm \lim_{x\to2^-}\color{orangered}{h(x)}=\lim_{x\to2^+}\color{orangered}{h(x)}\]We need these limits to agree.

zepdrix (zepdrix):

\[\Large\rm \lim_{x\to2^-}\color{orangered}{e^{kx}}=\lim_{x\to2^+}\color{orangered}{x+11}\]Do you understand how I was able to plug these in for h(x)?

OpenStudy (anonymous):

yes!

zepdrix (zepdrix):

Since we have no discontinuity (problem) at x=2, we can apply these limits by simply plugging in x=2 for each side.\[\Large\rm e^{2k}=2+11\]

zepdrix (zepdrix):

And then solve for k. Do you remember how to use your log rules? :O We'll need one of them here.

OpenStudy (anonymous):

ohh okay thank you! and i don't remember my log rules.. :( but i can search them up!

zepdrix (zepdrix):

This is the one that you'll want to remember:\[\Large\rm \ln(a^{\color{orangered}{b}})=\color{orangered}{b}\ln(a)\]

zepdrix (zepdrix):

Taking the log of each side will allow you to get the k OUT of the exponent position.

OpenStudy (anonymous):

ohh that's right! wow thank you.

OpenStudy (anonymous):

so when i log the 2 + 11 side, do i just calculate those values on my calculator?

zepdrix (zepdrix):

Mmmmm yah, or just leave it as a log. That probably is better actually. Hmm I'm trying to make sure I didn't make a boo boo somewhere, it's not graphing correctly for me.

zepdrix (zepdrix):

Ahhh nevermind it seems to be working out ok >.< I was just being silly I think.

zepdrix (zepdrix):

What'd you come up with?

OpenStudy (anonymous):

i haven't done the work yet! i'm working on another problem at the moment; once i'm finished i will post the answer here, though, if you'd like to check! thank you again

OpenStudy (anonymous):

i haven't simplified my work yet, but i got (ln2 + ln11)/2e ... i feel like this is wrong /:

OpenStudy (anonymous):

figured it out! i think! haha i got ln(13)/2

zepdrix (zepdrix):

Ah there we go \c:/ yay team!

OpenStudy (anonymous):

thank you!!